block A of mass 2 kg is kept on a block B of mass 3 kg. some friction exists between A &B but no friction between B and the ground. if 10N force on A produces 2m/s acceleration in it, then compute the acceleration in B

Grade:10

1 Answers

vikas askiitian expert
13 years ago

at upper block , friction(f) acts in backward ....

 F - f = mAaA = 2*2 = 4

 10 - f = 4

 f = 6N

at lower block , friction acts as driving force so

f = mBaB = 3a

6 = 3a

a = 2m/s2